How much work is done by the system when 2.00 moles of O2 expand from a volume of 3.00 liters to a volume of 9 liters against a constant external pressure of 0.0100 atm?
Answer in units of J
1
Expert's answer
2015-01-30T12:00:42-0500
W=P*change in vol
chnge in vol = 9 - 3 = 6 litres = 0.006m3
W=0.006*0.01*101325=6.0795 joule
Answer: 6.0795 joule